Which object has a greater momentum: a stationary tractor-trailer or a tossed tennis ball? Explain your answer.
Answer #1

Momentum is mass times velocity= m*v a tossed tennis ball has more momentum as it has some velocity and mass ,even though tractor trailer has huge mass it is stationary that means velocity is zero so momentum = m*0=0 kg m/sec.So tossed ball having some momentum has momentum greater than zero so tossed ball has greater momentum .